How many gallons of a 3% salt solution must be mixed with 30 gallons of a 9% solution to obtain a 6% solution?
Answer by mananth(16946)

percent ---------------- quantity
9 ---------------- 30
3 ---------------- x
6 ---------------- 30+x
...
9*30+3*x =6(30+x)
270+3x=180+6x
3x-6x=180-270
-3x=-90
/-3
x=30 liters